# TextRegion: Text-Aligned Region Tokens from Frozen Image-Text Models

### v1

@@ 4.2 Zero-shot Referring Expression Comprehension
Dataset. We evaluate on RefCOCO [44], RefCOCO+ [44] and RefCOCOg [21] datasets. All of these benchmarks come from the MS-COCO [18] dataset, paired with expressions that refer to a unique object in each image, accompanied with a bounding box. The main difference is that RefCOCO+ excludes relation-based expressions (e.g. “left of”, “closer”, or “bigger”) and RefCOCOg has longer descriptions. The test sets of RefCOCO and RefCOCO+ are divided into “testA” and “testB,” which contain only people and non-people instances, respectively.
Setting. Referring Expression Comprehension (ReC) focuses on identifying the most relevant object (bounding box) in an image based on a given natural language expression. A common zero-shot ReC setting leverages a pretrained object detector to generate object proposals and then selects the most likely proposal corresponding to the query. The accuracy is evaluated by the percentage of instances where the selected proposal has an Intersection-over-Union (IoU) of at least 0.5 with the ground-truth bounding box. For this experiment, we do not incorporate any spatial relationship refinement or post-processing methods commonly used in ReC. Retrieval accuracy is reported based on object proposals generated by MAttNet [45], and baseline results are taken from FGVP [42].
@@
CPT [43] VinVL [47] 32.3 36.1 30.3 31.9 35.2 28.8 36.7 36.5
TextRegion SigLIP2-L/16 [35] 45.0 49.8 36.0 51.3 57.1 40.5 53.8 52.6
TextRegion PE-Core-L/14 [1] 47.3 53.7 37.9 52.6 59.7 42.2 52.7 50.7
Red Circle [31] CLIP ViT-L/14@336 38.0 45.3 32.9 43.9 51.0 37.1 47.2 47.3
FGVP [42] CLIP ViT-L/14@336 46.1 53.0 40.4 50.4 57.5 42.6 54.5 54.1
@@
Results. Tab. 2 shows the results for zero-shot ReC. Complex baselines depend on carefully designed pipelines and prompting strategies [43, 31], and using bounding box candidates as prompts for SAM to obtain better object masks [42]. In contrast, our method directly generates region tokens without relying on candidate bounding boxes and still achieves superior performance.
@@ 4.3 Multiple Object Grounding
Setting. In Sec. 4.2 we test the referring expression comprehension performance of our method. But since that task involves referring to a single target object, we are interested in testing our method’s multi-object grounding capabilities. For this test, we use the Reasoning Segmentation dataset [13] which is commonly used to assess the visual segmentation capabilities of VLMs. As some samples in the dataset contain multiple objects related to a given query, simply returning the region with the highest similarity is not effective and can overlook relevant objects.
Since the original queries in the reasoning segmentation dataset are long and complex, we construct a simplified scenario. Specifically, we use the original LLaVA1.5-7B [19] to generate text answers for the original queries, and then treat these answers as interpreted queries to retrieve relevant objects. This scenario is designed to evaluate the impact of language understanding ability. Further details are provided in the supplementary material. We implement baseline results for Trident [29], while the results for all other baselines are sourced from LISA [13]. We use CLIP ViT-L/14@336px as the backbone for this experiment.
@@
OVSeg [17] 18.0 15.5 28.7 22.5 26.1 20.8
X-Decoder [50] 20.4 11.6 22.2 17.5 21.7 16.3
SEEM [51] 20.1 11.5 25.6 20.8 24.3 18.7
Grounded-SAM [20] 17.8 10.8 22.4 18.6 21.3 16.4
@@
OVSeg [17] 24.2 18.7 44.6 37.1 39.7 31.8
Trident [29] 23.0 24.2 43.3 42.5 38.4 39.2
@@
Metrics. There are two metrics for this task: gIoU and cIoU. gIoU is defined as the average of per-image Intersection-over-Unions (IoUs), while cIoU is computed as the cumulative intersection divided by the cumulative union. Since cIoU is heavily biased toward large-area objects and exhibits high variability, gIoU is usually preferred [13].
@@ 5 Ablation and Discussion
Global Patch, Multi-resolution and CLIP variants. We validate the effectiveness of the proposed global patch removal and multi-resolution patch feature strategies in Tab. 4. Overall, these strategies benefit a range of image-text models, from small to large backbones. When using both techniques, the performance is generally the best compared to using alone. An interesting observation is that the performance of SigLIP2-SO400M/16 drops when using the multi-resolution patch feature, whereas all other models benefit from it. A possible explanation is that large backbones, while powerful, tend to generate more global patches [41, 6, 28]. When applying the multi-resolution feature alone, this effect may amplified—simpler images contain less detailed information, causing more patches to capture global features. Our global patch removal strategy helps mitigate the negative effects, so when combined with multi-resolution patch features, it consistently delivers better or comparable performance than using multi-resolution alone.
Design choice for mask-based attention pooling. We evaluate different design choices for mask-based attention pooling using CLIP ViT-B/16. Tab. 5 compares the effect of different feature used for pooling. “Input” refers to the feature map fed into the final attention block, “Output” denotes the output of that block, and “Value” means the attention values from the last attention block. Interestingly, even with a single projection, transitioning from the input to the value features significantly improves performance—from almost zero to decent results. Tab. 6 shows the effect of pooling attention value from different
@@
layers. Results show that the value features from the second-to-last layer yields substantially better performance than directly pooling the input feature map (Tab. 5) for the last attention block. Tab. 7 shows the results of either upscaling the feature maps or downscaling the SAM2 masks when pooling region tokens. Experiments are conducted on COCO-Stuff. The reported “Up” result use a 5× upscaling of the value feature maps instead of directly matching the shape of the SAM2 masks, as the latter approach can induce much larger GPU consumption, particularly when combined with the multi-resolution strategy. “s/img” means seconds required to predict a single image, measured on one A100 GPU. The results show that downsampling offers both effectiveness and efficiency advantages. Tab. 8 shows the impact of using hard masks versus soft masks for aggregating region tokens and computing region logits.
@@ 6 Conclusion
Acknowledgement: This work is supported in part by ONR N00014-23-1-2383 and DARPA HR0011-23-9-0060. The views and conclusions expressed are those of the authors, and not necessarily representative of the US Government or its agencies.
Appendix A Supplementary materials
@@ A.1 Summary of contents
Sec. A.2 details the hyperparameters used in our experiments. In Sec. A.3, we describe how to generate text-aligned region tokens using the frozen SigLIP2 [35] and Perception Encoder [1]. Finally, Sec. A.4 explains the process of generating interpreted queries for Reasoning Segmentation.
@@ A.2 Hyperparameter
For all experiments, we filter global patches using a threshold of $\tau=0.07$. The crop size is uniformly set to 336 for all CLIP models (ViT-B/16 through ViT-H/14), while SigLIP2 and Perception Encoder use their respective default input resolutions. Region masks are generated with SAM2 [27] Hiera-Large, using the following configuration: pred-iou-thresh set to 0.6, stability-score-thresh to 0.6, box-nms-thresh to 0.9, and points-per-side to 16. In the semantic segmentation experiments on the Cityscapes dataset, we increase points-per-side to 36 to due to its high resolution and the abundance of small objects. To mitigate the impact of duplicated or overlapping masks, we also merge masks with an overlap IoU greater than 0.8.
@@ A.3 CLIP Variants
Unlike CLIP [25], which retains the [CLS] token across multiple layers, SigLIP2 [35] and Perception Encoder [1] introduce a learnable query token, denoted as [q], to aggregate global image information within the final attention pooling block. Let $X$ denote the input to this last attention pooling block:
@@
Another modification for SigLIP2 [35] and the Perception Encoder [1] is that we change the final value feature to $V_{\text{final}}=V_{\text{high}}+0.5*\text{upsample}\left(V_{\text{low}}\right)$ for the multi-resolution strategy, as they are more sensitive to the influence of global patches.
@@ A.4 Interpreted Queries for Multi-Object Grounding / Reasoning Segment
To convert the original long and complex queries in the reasoning segmentation dataset into interpretable queries compatible with the CLIP text encoder, we use LLaVA1.5-7B [19] for query preprocessing. Specifically, we input the paired image and query into LLaVA1.5-7B and obtain a concise answer using the prompt: “Please summarize the answer in five words or fewer to define the object.” The answer is then used as the interpreted query to obtain our multi-object grounding results. See Fig. 6 and Fig. 7 for more examples of interpreted queries.
@@
Figure 6: Examples for Interpreted Queries (Part 1): The title means the original query from the reasoning segmentation dataset, while the legend presents the interpreted query generated by LLaVA 1.5. The left column displays the query image. The middle column shows the ground truth, with green highlighting the target grounding area. The right column presents the grounding result.

### v2

@@ 4.2 Zero-shot Referring Expression Comprehension
Dataset. We evaluate on RefCOCO (yu2016modeling), RefCOCO+ (yu2016modeling) and RefCOCOg (mao2016generation) datasets. All of these benchmarks come from the MS-COCO (lin2014microsoft) dataset, paired with expressions that refer to a unique object in each image, accompanied with a bounding box. The main difference is that RefCOCO+ excludes relation-based expressions (e.g. “left of”, “closer”, or “bigger”) and RefCOCOg has longer descriptions. The test sets of RefCOCO and RefCOCO+ are divided into “testA” and “testB,” which contain only people and non-people instances, respectively.
Setting. Referring Expression Comprehension (ReC) focuses on identifying the most relevant object (bounding box) in an image based on a given natural language expression. A common zero-shot ReC setting leverages a pretrained object detector to generate object proposals and then selects the most likely proposal corresponding to the query. The accuracy is evaluated by the percentage of instances where the selected proposal has an Intersection-over-Union (IoU) of at least 0.5 with the ground-truth bounding box. For this experiment, we do not incorporate any spatial relationship refinement or post-processing methods commonly used in ReC. Retrieval accuracy is computed using object proposals produced by MAttNet (yu2018mattnet). All baseline results are taken from FGVP (yang2023fine). For a fair comparison, we evaluate TextRegion with the same set of object proposals as those used by FGVP.
@@
CPT (yao2024cpt) VinVL (zhang2021vinvl) 32.3 36.1 30.3 31.9 35.2 28.8 36.7 36.5
TextRegion SigLIP2-L/16 (tschannen2025siglip) 45.0 49.8 36.0 51.3 57.1 40.5 53.8 52.6
TextRegion PE-Core-L/14 (bolya2025perception) 47.3 53.7 37.9 52.6 59.7 42.2 52.7 50.7
Red Circle (shtedritski2023does) CLIP ViT-L/14@336 38.0 45.3 32.9 43.9 51.0 37.1 47.2 47.3
FGVP (yang2023fine) CLIP ViT-L/14@336 46.1 53.0 40.4 50.4 57.5 42.6 54.5 54.1
@@
Results. Tab. 2 shows the results for zero-shot ReC. Complex baselines depend on carefully designed pipelines and prompting strategies (yao2024cpt; shtedritski2023does), and using bounding box candidates as prompts for SAM to obtain better object masks (yang2023fine). In contrast, our method directly generates region tokens without relying on candidate bounding boxes and still achieves superior performance.
@@ 4.3 Multiple Object Grounding
OVSeg (liang2023open) 18.0 15.5 28.7 22.5 26.1 20.8
X-Decoder (zou2023generalized) 20.4 11.6 22.2 17.5 21.7 16.3
SEEM (zou2023segment) 20.1 11.5 25.6 20.8 24.3 18.7
Grounded-SAM (liu2024grounding) 17.8 10.8 22.4 18.6 21.3 16.4
@@
OVSeg (liang2023open) 24.2 18.7 44.6 37.1 39.7 31.8
Trident (shi2024harnessing) 23.0 24.2 43.3 42.5 38.4 39.2
@@
Setting. In Sec. 4.2 we test the referring expression comprehension performance of our method. But since that task involves referring to a single target object, we are interested in testing our method’s multi-object grounding capabilities. For this test, we use the Reasoning Segmentation dataset (lai2024lisa) which is commonly used to assess the visual segmentation capabilities of VLMs. As some samples in the dataset contain multiple objects related to a given query, simply returning the region with the highest similarity is not effective and can overlook relevant objects.
Since the original queries in the reasoning segmentation dataset are long and complex, we construct a simplified scenario. Specifically, we use the original LLaVA1.5-7B (liu2024improved) to generate text answers for the original queries, and then treat these answers as interpreted queries to retrieve relevant objects. This scenario is designed to evaluate the impact of language understanding ability. Further details are provided in Appendix Sec A.4. We implement baseline results for Trident (shi2024harnessing), while the results for all other baselines are sourced from LISA (lai2024lisa). We use CLIP ViT-L/14@336px as the backbone for this experiment.
Metrics. There are two metrics for this task: gIoU and cIoU. gIoU is defined as the average of per-image Intersection-over-Unions (IoUs), while cIoU is computed as the cumulative intersection divided by the cumulative union. Since cIoU is heavily biased toward large-area objects and exhibits high variability, gIoU is usually preferred (lai2024lisa).
@@ 5 Ablation and Discussion
Global Patch, Multi-resolution and CLIP variants. We validate the effectiveness of the proposed global patch removal and multi-resolution patch feature strategies in Tab. 4. Overall, these strategies benefit a range of image-text models, from small to large backbones. When using both techniques, the performance is generally the best compared to using alone. An interesting observation is that the performance of SigLIP2-SO400M/16 drops when using the multi-resolution patch feature, whereas all other models benefit from it. A possible explanation is that large backbones, while powerful, tend to generate more global patches (xiao2023efficient; darcet2023vision; shao2024explore). When applying the multi-resolution feature alone, this effect may amplified — simpler images contain less detailed information, causing more patches to capture global features. Our global patch removal strategy helps mitigate the negative effects, so when combined with multi-resolution patch features, it consistently delivers better or comparable performance than using multi-resolution alone.
Another interesting finding is that the global patch phenomenon is less severe in CLIP compared to the perception encoder and SigLIP2. As a result, using global patch removal alone for CLIP has a limited impact. However, when combined with multi-resolution features, it significantly boosts performance, particularly for CLIP ViT-H/14.
@@
Design choice for mask-based attention pooling. We evaluate different design choices for mask-based attention pooling using CLIP ViT-B/16. Tab. 5 compares the effect of different feature used for pooling. “Input” refers to the feature map fed into the final attention block, “Output” denotes the output of that block, and “Value” means the attention values from the last attention block. Interestingly, even with a single projection, transitioning from the input to the value features significantly improves performance—from almost zero to decent results.
Tab. 6 shows the effect of pooling attention value from different layers. Results show that the value features from the second-to-last layer yields substantially better performance than directly pooling the input feature map (Tab. 5) for the last attention block.
Tab. 7 shows the results of either upscaling the feature maps or downscaling the SAM2 masks when pooling region tokens. Experiments are conducted on COCO-Stuff. The reported “Up” result use a 5× upscaling of the value feature maps instead of directly matching the shape of the SAM2 masks, as the latter approach can induce much larger GPU consumption, particularly when combined with the multi-resolution strategy. “s/img” means seconds required to predict a single image, measured on one A100 GPU. The results show that downsampling offers both effectiveness and efficiency advantages. Tab. 8 shows the impact of using hard masks versus soft masks for aggregating region tokens and computing region logits.
Mask Generator COCO Stuff AED20K
Perception SigLIP2 CLIP Perception SigLIP2 CLIP
Ground Truth masks 34.7 33.7 36.5 32.7 35.4 31.0
SLIC (6205760) 21.7 19.9 21.3 18.4 18.2 16.1
SAM (kirillov2023segment) 27.0 26.1 27.9 24.3 25.8 22.1
SAM2 (ravi2024sam) 27.9 26.8 28.7 24.3 26.0 22.8
Table 9: Ablation on the effect of mask generators. “Ground Truth masks” refer to using the annotated masks to extract region tokens, which are subsequently used for prediction. All experiments are conducted with image-text model based on the ViT-B/16 backbone. The default configuration is highlighted in green.
Sensitivity to Mask Quality. Tab. 9 reports the impact of different mask generators. “Ground Truth Masks” means using the annotated ground-truth masks to compute region tokens, serving as the upper bound of TextRegion. In contrast, SLIC (6205760) partitions images into superpixels based on color and spatial cues, resulting in lower performance than SAM and SAM2. These results demonstrate that TextRegion is compatible with various segmentation methods, enabling object-level predictions for flexible, user-specified mask inputs. We also provide visual comparisons of those mask generators in Fig. 6.
Input Image Ground Truth GT Masks SLIC SAM SAM2
VOC21 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Context60
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
COCO Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Visualization comparison of different mask generators across datasets. “GT masks” refer to using the annotated masks to extract region tokens.
@@ 6 Conclusion
Acknowledgement: This work is supported in part by ONR N00014-23-1-2383 and DARPA HR0011-23-9-0060. This work used NVIDIA GPUs at NCSA Delta through allocation CIS240059 and CIS250059 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by NSF Grants #2138259, #2138286, #2138307, #2137603, and #2138296. The views and conclusions expressed are those of the authors, and not necessarily representative of the US Government or its agencies.
Appendix A Appendix
@@ A.1 Summary of contents
Sec. A.2 details the hyperparameters used in our experiments. In Sec. A.3, we describe how to generate text-aligned region tokens using the frozen SigLIP2 (tschannen2025siglip) and Perception Encoder (bolya2025perception). Sec. A.4 explains the process of generating interpreted queries for Reasoning Segmentation. Finally, Sec. A.5 includes further visual examples for open-world segmentation and referring expressions, as well as illustrative failure cases.
@@ A.2 Hyperparameter
For all experiments, we filter global patches using a threshold of $\tau=0.07$. The crop size is uniformly set to 336 for all CLIP models (ViT-B/16 through ViT-H/14), while SigLIP2 and Perception Encoder use their respective default input resolutions. Region masks are generated with SAM2 (ravi2024sam) Hiera-Large, using the following configuration: pred-iou-thresh set to 0.6, stability-score-thresh to 0.6, box-nms-thresh to 0.9, and points-per-side to 16. In the semantic segmentation experiments on the Cityscapes dataset, we increase points-per-side to 36 to due to its high resolution and the abundance of small objects. To mitigate the impact of duplicated or overlapping masks, we also merge masks with an overlap IoU greater than 0.8.
@@ A.3 CLIP Variants
Unlike CLIP (radford2021learning), which retains the [CLS] token across multiple layers, SigLIP2 (tschannen2025siglip) and Perception Encoder (bolya2025perception) introduce a learnable query token, denoted as [q], to aggregate global image information within the final attention pooling block. Let $X$ denote the input to this last attention pooling block:
@@
Another modification for SigLIP2 (tschannen2025siglip) and the Perception Encoder (bolya2025perception) is that we change the final value feature to $V_{\text{final}}=V_{\text{high}}+0.5*\text{upsample}\left(V_{\text{low}}\right)$ for the multi-resolution strategy, as they are more sensitive to the influence of global patches.
@@ A.4 Interpreted Queries for Multi-Object Grounding / Reasoning Segment
To convert the original long and complex queries in the reasoning segmentation dataset into interpretable queries compatible with the CLIP text encoder, we use LLaVA1.5-7B (liu2024improved) for query preprocessing. Specifically, we input the paired image and query into LLaVA1.5-7B and obtain a concise answer using the prompt: “Please summarize the answer in five words or fewer to define the object.” The answer is then used as the interpreted query to obtain our multi-object grounding results. See Fig. 7 and Fig. 8 for more examples of interpreted queries.
@@
Figure 7: Examples for Interpreted Queries (Part 1): The title means the original query from the reasoning segmentation dataset, while the legend presents the interpreted query generated by LLaVA 1.5. The left column displays the query image. The middle column shows the ground truth, with green highlighting the target grounding area, and red indicates areas that can be ignored. The right column presents the grounding result.
@@
A.5 Visualization
Input Image Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GT Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
TextRegion Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 9: Visualization for open-world segmentation. “GT” means the ground truth, while “TextRegion” shows our predicted results.
person holding umbrella dish in top right corner psn boy left guy child sitting on womans lap
GT Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
TextRegion Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 10: Visualization for referring expressions. The sentences above are the query expressions. The red bounding box in the “GT” row indicates the referred object.
Fig. 9 and Fig. 10 presents additional visualizations for open-world segmentation and referring expression tasks. From the visual results, we observe that while SAM2 can segment objects with relatively accurate boundaries, it struggles with objects exhibiting complex contours. For instance, in the train example of Fig. 9, TextRegion produces slight pixel-level inaccuracies along the intricate edges.
For referring expressions, the example of “person holding umbrella” shows that the text encoder may sometimes misinterpret the expression, leading the model to predict a related object rather than the specific target described.